OpenStudy (sambam):

How do solve this? (a^2 + 2a + 1)(a + 1)

OpenStudy (ankit042):

(a+b)^2=a^2+b^2+2ab

OpenStudy (ankit042):

can you identify value of b in the equation?

OpenStudy (anonymous):

Remember the formula a^2 + 2*a*b + b^2 = (a+b)^2 So , the b=1 a^2 +2a + 1=(a+1)^2 So , you can find the a : (a+1)^2=(a+1)(a+1) a=-1